# TrustGeoGen: Scalable and Formal-Verified Data Engine for Trustworthy Multi-modal Geometric Problem Solving

### v1

@@ 2.1 Data Engine
0:  Reasoning graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target state $s_{n}$, threshold $\tau_{l}$, threshold $\tau_{r}$
0:  Filtered reasoning path $\mathcal{P}$
1:  Initialize $\mathcal{P}\leftarrow\emptyset$
2:  $s\leftarrow s_{n}$
3:  while $s\notin S_{0}$ do
4:     Identify rule $r_{s}\in R$ and parent state set $S_{i-1}\subset S$ such that $S_{i-1}\xhookrightarrow{r_{s}}s$
5:     Append transition $(S_{i-1},r_{s},s)$ to $\mathcal{P}$
6:     $s\leftarrow S_{i-1}$
7:  end while
8:  Assess path length: $L\leftarrow|\mathcal{P}|$
9:  Compute premises ratio: $R_{P}\leftarrow|S_{P}|/|S_{0}|$, where $S_{P}\subseteq\mathcal{P}$
10:  if $L\geq\tau_{l}$ and $R_{P}\geq\tau_{r}$ then
11:     Return $\mathcal{P}$
12:  else
13:     Discard $\mathcal{P}$
14:  end if
Translator is responsible for converting the reasoning path $\mathcal{P}$ into a human-readable natural language explanation. To accomplish this, the Translator leverages a large language model (LLM) to ”smooth” the reasoning path. For each state and rule in the reasoning path $\mathcal{P}$, the Translator provides the LLM with a detailed natural language interpretation of the geometric states and inference rules, ensuring that the LLM fully understands the semantic and logical meaning of the components in $\mathcal{P}$. Additionally, the Translator utilizes a few-shot learning strategy to enhance translation accuracy and consistency. Specifically, it includes a pre-curated set of high-quality human-generated translations as few-shot examples, which serve to guide the LLM in generating precise and context-appropriate explanations. By iteratively processing each triplet $(S_{i-1},r_{s},s)\in\mathcal{P}$, the Translator generates a coherent chain-of-thought explanation that mirrors the logical flow of the reasoning graph $\mathcal{G}$. This ensures that the resulting translation not only reflects the exact logical derivations but also facilitates human understanding by articulating complex geometric relationships in an accessible and comprehensive manner.
@@
0:  Reasoning graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target state $s_{n}$, threshold $\tau_{l}$, threshold $\tau_{r}$, number of searches $n$
0:  Set of filtered reasoning paths $\mathcal{P}_{\text{set}}$
1:  Initialize $\mathcal{P}_{\text{set}}\leftarrow\emptyset$
[…]
3:  while True do
4:     Initialize $\mathcal{P}\leftarrow\emptyset$
5:     $s\leftarrow s_{n}$
6:     while $s\notin S_{0}$ do
7:        Identify set of rules $\mathcal{R}_{s}\subseteq R$ and corresponding parent state sets $\mathcal{S}^{m}_{i-1}\subseteq S$ such that $\forall r_{s}\in\mathcal{R}_{s},\exists S_{i-1}\in\mathcal{S}^{m}_{i-1}:S_{i-1}\xhookrightarrow{r_{s}}s$
8:        Select $r_{s}\in\mathcal{R}_{s}$ and $S_{i-1}\in\mathcal{S}^{m}_{i-1}$ (choose different options in different searches)
[…]
10:        if $(S_{i-1},r_{s},s)\notin\mathcal{P}$ then
11:           $\mathcal{P}\leftarrow\mathcal{P}\cup(S_{i-1},r_{s},s)$
12:        end if
13:        $s\leftarrow S_{i-1}$
14:     end while
15:     Assess path length: $L\leftarrow|\mathcal{P}|$
16:     Compute premises ratio: $R_{P}\leftarrow|S_{P}|/|S_{0}|$, where $S_{P}\subseteq\mathcal{P}$
17:     if $L\geq\tau_{l}$ and $R_{P}\geq\tau_{r}$ then
18:        $\mathcal{P}_{\text{set}}\leftarrow\mathcal{P}_{\text{set}}\cup\{\mathcal{P}\}$
19:     end if
20:     if used_options contains all possible options then
21:        Break
22:     end if
23:  end while
24:  return  $\mathcal{P}_{\text{set}}$
@@ 2.3 Multi-solution Data
[…]
@@
0:  Directed graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target state $s_{t}\in S$, threshold $\tau_{p}$
0:  Trace-back reasoning data $\mathcal{D}$
1:  Initialize $\mathcal{D}\leftarrow\emptyset$
2:  Compute all reasoning paths $\mathcal{P}^{t}_{\text{set}}=\{\mathcal{P}_{1}^{t},\mathcal{P}_{2}^{t},\dots\}$ to $s_{t}$ using Algorithm 2
3:  Randomly sample a state $s_{e}\in S$ such that $s_{e}\notin\text{UpstreamDependencies}(\mathcal{P}^{t})$ for any $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$
4:  Search for a reasoning path $\mathcal{P}^{e}$ leading to $s_{e}$
5:  if $\exists\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}},\text{Overlap}(\mathcal{P}^{e},\mathcal{P}^{t})\geq\tau_{p}$ then
6:     Add $\mathcal{P}^{e}\cup\mathcal{P}^{t}$ to $\mathcal{D}$
7:  end if
8:  return  $\mathcal{D}$
@@ 2.4 Self-reflective Traceback Data
Recent studies on the reasoning capabilities of large models suggest that these models acquire substantial knowledge during the pre-training phase. By providing training data with various cognitive templates during the post-training phase, the models can better utilize their inherent knowledge, thereby enhancing reasoning ability and improving response quality[42, 41]. Among these cognitive templates, self-reflection trace-back reasoning serve as highly effective approaches. However, self-reflection and retrospective reasoning often occur during the author’s cognitive process and are rarely preserved in the final outputs, making the collection of relevant data particularly challenging. These types of data are even more scarce in the domain of multi-modal geometric problem-solving.
TrustGeoGen defines trace-back reasoning data as a sub-graph of $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, characterized as the union of sub-graphs with identical upstream dependencies but differing final states. Specifically, for a given piece of trace-back reasoning data, the process involves first deriving an incorrect final state, then retrospectively navigating a correct reasoning path (shared by both the incorrect and correct paths), and ultimately arriving at the correct final state. To achieve this, TrustGeoGen selects a target state $s_{t}$ and employs Algorithm 2 to enumerate all possible reasoning paths $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$. Next, a state $s_{e}$ is randomly sampled from the state set $S$, ensuring that $s_{e}$ does not belong to the upstream dependencies of any $\mathcal{P}^{t}$. Subsequently, TrustGeoGen searches for a reasoning path $\mathcal{P}^{e}$ leading to $s_{e}$. If $\mathcal{P}^{e}$ overlaps with any $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$ and the overlapping portions exceed the threshold $\tau_{p}$, TrustGeoGen outputs $\mathcal{P}^{e}\cup\mathcal{P}^{t}$ as trace-back reasoning data. Notably, Algorithm 2 is utilized to identify all possible reasoning paths for $s_{t}$ to ensure that $s_{e}$ does not appear in any potential upstream dependencies, thereby maintaining the integrity of the reasoning process leading to $s_{t}$.The details of the trace-back data acquisition algorithm, GeoExplore-T, are presented in Algorithm 3.
@@
Figure 4: Visualization examples of different difficulty levels in GeoTrust-test, where ”reasoning length” indicates step length of reasoning path and ”premise ratio” refer to the ratio of premises unutilized during formal reasoning and premises provided in the question.
@@ 3 Data Analysis
The modality-completed trustworthy geometric reasoning data holds great potential for improving the geometric problem-solving capabilities of MLLMs. Moreover, it could generalize these abilities to other tasks requiring deep reasoning. To validate it, we utilized TrustGeoGen to produce the dataset GeoTrust of 200K samples, from which 8K were sampled as the training set. Additionally, we employed different thresholds, $\tau_{l}$ and $\tau_{r}$, in conjunction with manual screening to curate a testset, GeoTrust-test, with 240 samples of varying levels of difficulty.
@@ 3.1 Data Distribution
TrustGeoGen was executed on a 60-core Intel Xeon (32) CPU @ 2.900GHz for two days, generating an initial dataset of 200K geometric reasoning instances in formal language. During the data generation process, we set the two filtering thresholds in the Sampler to $\tau_{l}=5$ and $\tau_{r}=0.5$, respectively. The distributions of the reasoning length and premises ratio for the resulting 200K dataset are illustrated in Fig. 3(a) and Fig. 3(b), respectively. The majority of the data is concentrated in the region where the reasoning length is less than 60, and it decreases sharply beyond 80. However, a small portion of the data exhibits reasoning lengths exceeding 150, indicating a considerably high level of reasoning complexity. For a more detailed view of the distribution, please refer to the zoomed-in section of Fig. 3(a).
Table 1: Performance of state-of-the-art multi-modal language models on GeoTrust-test, which is divided into four difficulty levels.
@@
num acc num acc num acc num acc num acc
LLaVA-1.5 [19] 7B 7 2.91% 3 5.00% 2 3.33% 0 0.00% 2 3.33%
Qwen2-VL-Instruct [36] 7B 9 4.17% 5 8.33% 2 3.33% 0 0.00% 2 3.33%
Qwen2.5-VL-Instruct [5] 7B 24 10.00% 12 20.00% 6 10.00% 3 5.00% 3 5.00%
GPT-4o [24] - 55 22.92% 25 41.67% 15 25.00% 9 15.00% 6 10.00%
Claude-3-7-Sonnet [3] - 83 34.58% 35 58.33% 17 28.33% 24 40.00% 7 11.67%
DeepSeek-R1 [10] 671B 108 45.00% 40 66.67% 24 40.00% 25 41.67% 19 31.67%
OpenAI-o1 [25] - 118 49.17% 42 70.00% 28 46.67% 26 43.33% 22 36.67%

### v2

@@ 20 Discard $\mathcal{P}$;
Reasoner leverages a predefined set of geometric theorems to infer new statements from premises generated by the constructor, formulating a reasoning graph:
$\small\mathcal{G}=(S,S_{0},R,\hookrightarrow),$ (1)
  • $S$ is the set of statements, where each statement $s\in S$ represents a derived conclusion or fact in the reasoning process.
  • $S_{0}\subseteq S$ is the set of initial statements generated by the constructor.
  • $R$ is the set of rules, where each rule $r\in R$ defines a logical inference relationship between statements.
  • $\hookrightarrow\ \subseteq S\times R\times S$ = $\{(S_{r},r,s^{\prime})\mid S_{r}\subset S,r\in R,s^{\prime}\in S\}$ is the statement transition relation. A transition $S_{r}\xhookrightarrow{r}s^{\prime}$ indicates statement $s^{\prime}$ is derived from $s$ by applying rule $r$.
Reasoning begins with the initial statement $S_{0}$ (given premises from the constructor), forming the root nodes of the graph $\mathcal{G}$. The graph expands by applying inference rules to existing statements. For every statement set $S_{r}\subset S$ and rule $r\in R$, if $r$ can be applied to $S_{r}$ to derive a new statement $s^{\prime}$ (formally $S_{r}\xhookrightarrow{r}s^{\prime}$), the reasoner performs two atomic updates: 1) Add the new statement $S_{r}$ to $S$ and 2) Add the edges $(S_{r},s^{\prime})$ to the transition relation $\hookrightarrow$ with rule $r$. Finally, the complete graph is the smallest closure satisfying:
$\small S=S_{0}\cup\left\{s^{\prime}\mid\forall S_{r}\subset S,\,r\in R,\,S_{r}\xhookrightarrow{r}s^{\prime}\right\}.$ (2)
Sampler operates on the reasoning graph $\mathcal{G}$ by sampling a target statement $s_{n}$ and subsequently searching for its path:
$\small\mathcal{P}=\{(S_{i-1},r_{s},s)\mid\forall s\in S_{i},S_{i-1}\xhookrightarrow{r_{s}}s,i=n,\dots,1\},$ (3)
where each triplet $(S_{i},r_{s},s)$ denotes a specific transition during the reasoning process, derived by applying rule $r_{s}$ to statement set $S_{i-1}$ to produce statement $s$. Any statement $s_{n}$ within the statement set $S$ can be treated as a problem to solve in order to derive a reasoning path $\mathcal{P}$. The sampler is initialized with a statement $s\in S_{n}$ to retrieve the corresponding transition rule $r_{s}$ and the upstream dependent statements $S_{i}$. This process is repeated iteratively until the statement set $S_{i}$ consists entirely of initial premises $S_{0}$. At that point, the full reasoning path has been successfully constructed. Notably, during the graph construction process in the Reasoner, once a statement $s$ is derived through rule $r_{s}$, no additional statement transition relations are appended to $s$. Consequently, the Sampler is always able to identify a unique inference rule $r_{s}$ corresponding to any statement $s$, thereby guaranteeing the absence of cycles in the reasoning path.
To ensure the quality of the reasoning paths, the Sampler employs two metrics to filter out unsuitable paths. The first metric evaluates the length of the reasoning path $\mathcal{P}$ against a predefined threshold $\tau_{l}$. If the length of $\mathcal{P}$ exceeds $\tau_{l}$, the path is considered difficult enough. Formally, this condition is expressed as $|\mathcal{P}|\geq\tau_{l}$, where $|\mathcal{P}|$ denotes the number of transitions in the reasoning path $\mathcal{P}$. The second metric assesses the ratio of the statements in the reasoning path $\mathcal{P}$ that are derived from the initial premises set $S_{0}$. If this ratio exceeds a predefined threshold $\tau_{r}$, the path is deemed sufficient enough. This condition is formally defined as $\frac{|S_{P}|}{|S_{0}|}\geq\tau_{r}$, where $|S_{P}|$ represents the number of initial premises used in the reasoning path $\mathcal{P}$, and $|S_{0}|$ is the number of all initial premises. By applying these two metrics, the Sampler ensures that only high-quality reasoning paths are retained, thereby maintaining the integrity and usefulness of the generated data. The detailed procedure of the aforementioned reasoning path exploration and filtering algorithm, GeoExplore, is presented in  algorithm˜1.
Translator is tasked with converting the reasoning steps synthesized by a formal engine into natural language narratives. These narratives serve as high-quality training data to cultivate the reasoning capabilities of Large Language Models (LLMs). As illustrated in  fig.˜2, the Translator’s process is twofold. First, it employs a state-of-the-art LLM (GPT-4o) to translate each formal reasoning step into its natural language counterpart. Given the highly structured format of the formal steps, we utilize a few-shot prompting strategy. This approach facilitates a step-by-step translation process rather than a single-pass generation of the entire sequence, thereby significantly enhancing translation fidelity.
However, a direct translation, while lexically aligned with human language, still mirrors the discrete, logical leaps of the formal engine, differing substantially from human cognitive processes. A key limitation is the absence of explicit logical bridges between consecutive steps, failing to articulate how one deduction leads to the next. Training an LLM on such data risks encouraging it to overfit to the syntactic structure of the formal expressions, rather than internalizing the underlying reasoning process. To mitigate this, the Translator executes a second crucial task: interleaving bridging rationales. It uses the translated steps as a scaffold and prompts GPT-4o to articulate the intermediate thought process. Specifically, for each step, the model first summarizes the currently established facts, then explicates how these facts logically connect to the subsequent step, and how this progression contributes to reaching the final goal. This method of generating explicit "Connection Thinking" compels the model to engage in a deeper level of inference, ultimately enhancing the reasoning performance of the models trained on this data.
Figure 2: Simple illustration of how Translator works
1
2
@@
Require: Reasoning graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target statement $s_{n}$, threshold $\tau_{l}$, threshold $\tau_{r}$, number of searches $n$
Ensure: Set of filtered reasoning paths $\mathcal{P}_{\text{set}}$
3
4Initialize $\mathcal{P}_{\text{set}}\leftarrow\emptyset$;
5 Initialize $\text{used\_options}\leftarrow\emptyset$;
6
7while True do
8 Initialize $\mathcal{P}\leftarrow\emptyset$;
9 $s\leftarrow s_{n}$;
10
11 while $s\notin S_{0}$ do
12      Identify set of rules $\mathcal{R}_{s}\subseteq R$ and corresponding parent statement sets $\mathcal{S}^{m}_{i-1}\subseteq S$ such that $\forall r_{s}\in\mathcal{R}_{s},\exists S_{i-1}\in\mathcal{S}^{m}_{i-1}:S_{i-1}\xhookrightarrow{r_{s}}s$;
13      Select $r_{s}\in\mathcal{R}_{s}$ and $S_{i-1}\in\mathcal{S}^{m}_{i-1}$ (choose different options in different searches);
14    $\text{used\_options}\leftarrow\text{used\_options}\cup\{(r_{s},S_{i-1})\}$;
15
16    if $(S_{i-1},r_{s},s)\notin\mathcal{P}$ then
17       $\mathcal{P}\leftarrow\mathcal{P}\cup(S_{i-1},r_{s},s)$;
18
19    $s\leftarrow S_{i-1}$;
20
21
22  Assess path length: $L\leftarrow|\mathcal{P}|$;
23   Compute premises ratio: $R_{P}\leftarrow|S_{P}|/|S_{0}|$, where $S_{P}\subseteq\mathcal{P}$;
24
25 if $L\geq\tau_{l}$ and $R_{P}\geq\tau_{r}$ then
26    $\mathcal{P}_{\text{set}}\leftarrow\mathcal{P}_{\text{set}}\cup\{\mathcal{P}\}$;
27
28
29 if used_options contains all possible options then
30    Break;
31
32
33return $\mathcal{P}_{\text{set}}$;
@@ 3.3 Multi-solution Data
During the graph construction phase, in order to obtain multiple solutions, for each statement $s$, when the Reasoner repeatedly identifies transitions in the form $S_{r}\xrightarrow{r}s$, each identified transition is retained. This retention of multiple transitions provides potential paths for exploring diverse solutions. In the solving process, the Sampler employs an adaptive search algorithm to generate reasoning paths. For any given problem $s_{n}$, the Sampler performs multiple path searches. If a statement $s_{i}$ has multiple outgoing transitions, the Sampler selects a different transition at each iteration for the next reasoning step. Consequently, the reasoning paths obtained during these iterations will differ. It is worth noting that to ensure the generated reasoning paths are acyclic, the Sampler enforces a constraint during path construction. Specifically, when attempting to add a triplet $(S_{i},r_{s},s)$ to the reasoning path $\mathcal{P}$, the Sampler checks whether this triplet already exists in $\mathcal{P}$. If it does, the statement $s$ is not revisited, thereby preventing cyclical paths in the reasoning process. For details of the multi-solution path searching algorithm, GeoExplore-M, refer to  algorithm˜2.
1
2
Require: Directed graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target statement $s_{t}\in S$, threshold $\tau_{p}$
Ensure: Trace-back reasoning data $\mathcal{D}$
3
4Initialize $\mathcal{D}\leftarrow\emptyset$;
5 Compute all reasoning paths $\mathcal{P}^{t}_{\text{set}}=\{\mathcal{P}_{1}^{t},\mathcal{P}_{2}^{t},\dots\}$ to $s_{t}$ using algorithm˜2;
6 Randomly sample a statement $s_{e}\in S$ such that $s_{e}\notin\text{UpstreamDependencies}(\mathcal{P}^{t})$ for any $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$;
7 Search for a reasoning path $\mathcal{P}^{e}$ leading to $s_{e}$;
8
9if $\exists\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}},\text{Overlap}(\mathcal{P}^{e},\mathcal{P}^{t})\geq\tau_{p}$ then
10   Add $\mathcal{P}^{e}\cup\mathcal{P}^{t}$ to $\mathcal{D}$;
11
12
13return $\mathcal{D}$;
@@ 3.4 Self-reflective Traceback Data
Figure 5: Visualization examples of different difficulty levels in GeoTrust-test, where "reasoning length" indicates step length of reasoning path and "premise ratio" refer to the ratio of premises unutilized during formal reasoning and premises provided in the question.
Recent studies on the reasoning capabilities of large models suggest that these models acquire substantial knowledge during the pre-training phase. By providing training data with various cognitive templates during the post-training phase, the models can better utilize their inherent knowledge, thereby enhancing reasoning ability and improving response quality(ye2025limo, ; yang2025reasonflux, ). Among these cognitive templates, self-reflection trace-back reasoning serve as highly effective approaches. However, self-reflection and retrospective reasoning often occur during the author’s cognitive process and are rarely preserved in the final outputs, making the collection of relevant data particularly challenging. These types of data are even more scarce in the domain of multi-modal geometric problem-solving.
TrustGeoGen defines trace-back reasoning data as a sub-graph of $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, characterized as the union of sub-graphs with identical upstream dependencies but differing final statement. Specifically, for a given piece of trace-back reasoning data, the process involves first deriving an incorrect final statement, then retrospectively navigating a correct reasoning path (shared by both the incorrect and correct paths), and ultimately arriving at the correct final statement. To achieve this, TrustGeoGen selects a target statement $s_{t}$ and employs algorithm˜2 to enumerate all possible reasoning paths $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$. Next, a statement $s_{e}$ is randomly sampled from the statement set $S$, ensuring that $s_{e}$ does not belong to the upstream dependencies of any $\mathcal{P}^{t}$. Subsequently, TrustGeoGen searches for a reasoning path $\mathcal{P}^{e}$ leading to $s_{e}$. If $\mathcal{P}^{e}$ overlaps with any $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$ and the overlapping portions exceed the threshold $\tau_{p}$, TrustGeoGen outputs $\mathcal{P}^{e}\cup\mathcal{P}^{t}$ as trace-back reasoning data. Notably, algorithm˜2 is utilized to identify all possible reasoning paths for $s_{t}$ to ensure that $s_{e}$ does not appear in any potential upstream dependencies, maintaining the integrity of the reasoning process leading to $s_{t}$. The details of the trace-back data acquisition algorithm, GeoExplore-T, are presented in  algorithm˜3.
@@ 4 Data Analysis
The modality-completed trustworthy geometric reasoning data holds great potential for improving the geometric problem-solving capabilities of MLLMs. Moreover, it could generalize these abilities to other tasks requiring deep reasoning. To validate it, we utilized TrustGeoGen to produce the dataset GeoTrust of 200K raw samples [^3] Raw samples refers to the data which is not processed by translator, from which 8K were sampled as the training set. Additionally, we employed different thresholds, $\tau_{l}$ and $\tau_{r}$, in conjunction with manual screening to curate a testset, GeoTrust-test, with 240 samples of varying levels of difficulty.
@@ 4.1 Data Distribution
TrustGeoGen was executed on a 60-core Intel Xeon (32) CPU @ 2.900GHz for two days, generating an initial dataset of 200K geometric reasoning instances in formal language. During the data generation process, we set the two filtering thresholds in the Sampler to $\tau_{l}=5$ and $\tau_{r}=0.5$, respectively. The distributions of the reasoning length and premises ratio for the resulting 200K data are illustrated in  fig.˜4(a) and  fig.˜4(b), respectively. The majority of the data is concentrated in the region where the reasoning length is less than 60, and it decreases sharply beyond 80. However, a small portion of the data exhibits reasoning lengths exceeding 150, indicating a considerably high level of reasoning complexity. For a more detailed view of the distribution, please refer to the zoomed-in section of  fig.˜4(a).
TABLE II: Performance of state-of-the-art multi-modal language models on GeoTrust-test, which is divided into four difficulty levels. The best results are highlighted in bold, and the second best results are highlighted with an underline.
@@
Count Accuracy Count Accuracy Count Accuracy Count Accuracy Count Accuracy
LLaVA-1.5-7B (llava/liu2024visual, ) 7B 4 1.67% 3 5.00% 1 0.42% 0 0.00% 0 0.00%
Qwen2-VL-7B (Qwen2-VL, ) 7B 11 4.58% 5 8.33% 2 3.33% 2 3.33% 2 3.33%
GPT-4o openai2024gpt4o - 62 25.83% 31 51.67% 10 16.67% 11 18.33% 10 16.67%
Claude-3.7-sonnet (2024claude, ) - 66 27.50% 33 55.00% 16 26.67% 10 16.67% 7 11.67%
Qwen2.5-VL-72B Qwen2.5-VL 72B 68 28.33% 32 53.33% 15 25.00% 15 25.00% 6 10.00%
DeepSeek-R1 * (DeepSeekAI2025DeepSeekR1IR, ) 671B 89 37.08% 30 50.00% 20 33.33% 22 36.67% 17 28.33%
Intern-S1 bai2025intern 235B+6B 104 43.33% 35 58.33% 21 35.00% 25 41.67% 23 38.33%
Gemini-2.5-pro comanici2025gemini - 104 43.33% 34 56.67% 24 40.00% 26 43.33% 20 33.33%
OpenAI-o3 openai_o3 - 110 45.83% 37 61.67% 24 40.00% 26 43.33% 23 38.33%
* For DeepSeek-R1, we only provide natural language questions, but no images.